# The immersed boundary method: an accelerated SIMPLE approach for moving bodies

### v1

@@ Abstract
We present a novel and scalable approach to accelerate a direct-forcing IBM for moving boundary simulations. Building upon the SIMPLE approach, we keep the pressure and forces coupled, formulate the pressure-forces corrections system, solve for pressure-forces corrections, then perform a projection-correction step for velocities followed by a correction step for pressure and forces. To solve the pressure-forces system, we apply a block reduction and suggest an efficient preconditioning approach that is based on using the Laplacian as a preconditioner for the resulting Schur complement. The Laplacian, being independent of the coordinates of the immersed body, is factorized only once (using an efficient direct method) and is then reused throughout the simulation. We present a rigorous proof of the spectral equivalence of the Laplacian and the Schur complement, thereby showing that the preconditioner’s efficiency is independent of the grid resolution.
The performance of the method is validated through simulations of the flow around a periodically oscillating sphere in a three-dimensional box. The methodology is then applied to analyze flows generated by arrays of 7 and 14 sub-spheres modeling porous spheres of different porosities. We show good scalability of the algorithm in terms of computational time and memory consumptions. That is, the iteration count for the preconditioned pressure-forces system is nearly constant, regardless of the grid resolution, Reynolds number, and the number of spheres. The developed approach enables accurate moving boundary simulations to be performed on standard workstations, thereby enhancing accessibility of the simulations to the CFD community.
@@
Moving boundary simulations, Implicit immersed boundary method, Schur complement, Preconditioning
@@ 1 Introduction
The immersed boundary method (IBM), introduced about 50 years ago by Peskin [1], remains a popular approach in investigations of fluid-structure interactions (FSI) for handling boundary conditions at the interfaces of bodies with complex geometries immersed in a fluid. The IBM is applicable to a wide range of physical and biological problems, including flow around solid bodies or inside elastic compartments, and particulate and two-phase immiscible flows, to name but a few. For a comprehensive overview, see the recent review [2] and references therein.
Among all the IBM variants, the direct-forcing approach [3, 4] has gained prominence in recent decades. Its popularity stems from the simplicity of mesh formation and the ability to directly impose no-slip constraints, without the need for assumptions regarding system dynamics. In the direct forcing approach, flow variables are discretized on a regular (Eulerian) grid, while the forces applied by the body on the surrounding flow are discretized on a curved (Lagrangian) grid attached to the body surface. Closure of the overall system is achieved by adding no-slip kinematic constraints. Since the Eulerian and Lagrangian grids do not necessarily coincide, there is a need to define two adjoint operators—interpolation and regularization—to enable information transfer between the two grids. The interpolation operator transfers velocities from the Eulerian to the Lagrangian grid, while the regularization operator smears the Lagrangian body forces over the adjacent Eulerian grid.
Many extensions of the original direct-forcing IBM have been suggested over the years, e.g, for obtaining high-order accuracy in satisfying no-slip constraints [5, 6, 7, 8] for better fulfilling the divergence-free condition [9], for application of the method on unbounded domains [10], and for enforcing the immersed boundary conditions only on one side of the immersed boundary by using Heaviside functions [11]. However, even in the original low-order IBMs, there is a need for a fast method to solve the resulting discrete system, especially for large 3D configurations hosting moving bodies: since the velocities, pressure, and kinematic and dynamic constraints in such a system are coupled at each time step, the system is very large and ill-conditioned, and a monolithic solution is not practical. A variety of approaches have been suggested to address this problem; these include fully explicit [1], fully implicit, and semi-implicit approaches [12, 13, 14, 15, 16], and projection methods [17, 18].
Among these approaches, the immersed boundary projection method [17] keeps the pressure and forces coupled and applies a projection method to the velocity field. An intermediate velocity field is used to construct a modified Poisson equation on the pressure and forces, and a projection step is then applied on the velocity to enforce the divergence-free constraints. To accelerate the solution, the modified Poisson equation is formed by using a Taylor expansion in time rather than a direct inverse of the leading block. However, the modified Poisson equation remains ill-conditioned, becoming particularly restrictive for 3D problems with a high grid resolution. Furthermore, efficient methods available for solving standard Poisson equations on structured grids cannot be fully leveraged for this formulation. In [15], an alternative approach is proposed, employing a Schur complement reduction to implicitly couple the non-solenoidal velocity and Lagrangian force fields, followed by a standard projection-correction procedure to obtain the correct pressure and divergence-free velocity fields. Unfortunately, the method requires a very long pre-computing time (about two weeks) for moving bodies, and extensive storage space for the pre-computed LU factors of the Schur complements that varies for each time step.
Recently, the SIMPLE-based approach for IBM was suggested in [16]. There, an intermediate velocity field is first calculated by solving the momentum equation. Thereafter, a generalized saddle-point system is formulated, consisting of a Poisson equation for the pressure corrections field coupled with an equation on the forces corrections field, that enforces the kinematic no-slip constraints on the surface of the immersed body. This saddle-point matrix is equivalent to the matrix that appears in the modified Poisson equation in [17], for the case of a first-order Taylor approximation. The pressure-forces corrections system is solved in [16] by a block reduction which yields the forces corrections field, from which the pressure corrections field is then obtained. The main drawback of this method is that it is prohibitively expansive when applied to moving-boundary configurations, since the Schur complement which constitutes the force corrections equation changes at each time step.
In this work, we adapt the SIMPLE approach [16] to accommodate moving boundary configurations. We achieve a significant reduction in computational time and a notably low memory usage. The pressure-forces corrections system—whose solution constitutes the computational bottleneck—is solved by taking the primal Schur complement, rather than the dual Schur complement used in [16], i.e., obtaining the pressure corrections field first, from which the force-corrections field is then computed. In this way, the overall complexity of the solution is significantly reduced by virtue of the approximation of the regularization block by a scalar matrix, as suggested in [16]. To further improve the conditioning and to enhance the scalability of the overall method through the solution of the pressure-corrections equation, we have developed a novel preconditioning approach, in which the Laplacian is taken as a preconditioner for the primal Schur complement of the pressure-forces corrections system. In other words, we solve the pressure corrections equation by using a preconditioned Krylov method, in which each iteration requires the solution of a single Poisson equation. Thanks to the special structure of the Laplacian on a regular grid, the Poisson equation is solved by a fast direct method [19, 20]. Moreover, since the Laplacian matrix is constant throughout the time steps, it needs to be factorized only once at the beginning of the computation.
Overall, our accelerated SIMPLE-based IBM shows promising scaling properties in terms of memory consumption and computational time. We prove rigorously and demonstrate numerically that the conditioning properties of the suggested preconditioning approach does not depend on the grid size or other system parameters. We demonstrate the potential advantages of our approach by applying it to the numerical simulation of the flow developing around a periodically oscillating sphere immersed within a 3D box. The methodology is further utilized to investigate more complex flows generated by arrays of 7 and 14 sub-spheres modeling porous spheres with different porosities. We thereby show that our method is suitable for performing realistic moving boundary simulations on standard workstations.
2 Background
2.1 Governing equations and discretization
@@
The incompressible flow around an immersed body is governed by the non-dimensional Navier-Stokes equations, complemented by no-slip constraints on the body’s surface:
$\frac{\partial{\bf u}}{\partial t}+({\bf u}\cdot\nabla){\bf u}=-\nabla p+\frac{1}{Re}\nabla^{2}{\bf u}+{\bf f}$ $\forall\vec{x}\in\Omega,$ (1)
$\nabla\cdot{\bf u}=0$ $\forall\vec{x}\in\Omega,$ (2)
${\bf u}={\bf U}^{\Gamma}$ $\forall\vec{x}\in\Gamma,$ (3)
BC $\forall\vec{x}\in\partial\Omega,$ (4)
where $\Omega$ is a domain, $\Gamma\subset\Omega$ is the boundary of the immersed body (or bodies), $p=p(\vec{x},t)$ is the pressure, ${\bf f}={\bf f}(\vec{x},t)$ reflects the impact of immersed body on surrounding flow, ${\bf u}$ is the flow velocity, and ${\bf U}^{\Gamma}$ is the velocity of the surface of the immersed body. The normalized equations (1), (2) and (3) are governed by a single non-dimensional parameter: the Reynolds number, $Re=U_{0}L/\nu$ where $\nu$ is the kinematic viscosity and $L$ and $U_{0}$ are the characteristic length and velocity, respectively. The time, pressure and force density are normalized by utilizing the scales $L/U_{0}$, $\rho U_{0}^{2}$ and $\rho U_{0}^{2}/L$, respectively, where $\rho$ is the fluid density.
The flow velocity and pressure, constituting the Eulerian variables, are discretized on a staggered grid: pressure is stored at cell centers, while velocity components are located at the corresponding face centers. The spatial derivatives are discretized using a standard central second-order finite volumes scheme. The time derivative is discretized by a backward second order finite differences scheme.
Following the IBM formalism, the surface of immersed body is determined by a series of discrete points, constituting the Lagrangian grid, see Fig. 1. For numerical stability, the distance between adjacent Lagrangian points should be approximately equal to the Eulerian cell size. Variables on the Lagrangian grid are denoted by capital letters, e.g., ${\bf F}$ corresponds to the Lagrangian forces discretized at the Lagrangian points.
@@
Since the Eulerian and Lagrangian grids do not necessarily coincide, two adjoint operators are next introduced to facilitate the data transfer between them: the regularization operator $R$, which smears the Lagrangian forces onto the underlying Eulerian grid, and the interpolation operator $R^{T}$, which interpolates the Eulerian velocities to Lagrangian points. These operators are defined by utilizing the following discrete Dirac delta function, proposed in [21]:
@@
where $r$ represents the distance from the smeared point and $\Delta r$ represents the cell width in the $r$ direction. This delta function gained popularity over the years [22, 23, 17, 16], due to its compact support of at most three discretization cells in each direction. [^1] Since the discrete Dirac delta function in Eq. (5) provides only a first order of spatial accuracy, other delta functions can be also considered to improve the accuracy (with the price of less favorable sparsity), see [6].
The regularization and interpolation operators are then defined, using the discrete Dirac delta function from Eq. (5), by
${\bf U}_{k}^{\Gamma}=(R^{T}{\bf u})_{k}=\Delta x^{3}\sum_{i}{\bf u}_{i}\delta^{3}({\bf X}_{k}-{\bf x}_{i}),$ (6)
${\bf f}_{i}=(R{\bf F})_{i}=\Delta x^{3}\sum_{k}{\bf F}_{k}\delta^{3}({\bf x}_{i}-{\bf X}_{k}),$ (7)
where ${\bf x}_{i}$ are the coordinates of the $i$-th velocity location, and ${\bf X}_{k}$ is the $k$-th Lagrangian point. The notation $\delta^{3}$ corresponds to the discrete Dirac delta function applied in 3D. As long as we use the same delta function for the interpolation and the regularization, the two operators are transpose of each other in matrix representation. To facilitate the notation, we denote by $R$ the regularization and by $R^{T}$ the interpolation operator.
2.2 The SIMPLE approach for IBM
We recall that the SIMPLE algorithm [24] for solution of the Navier-Stokes equations starts with a prediction step for the calculation of an intermediate velocity field ${\bf u}^{*}$, which is subsequently used in the calculation of the corrections for the pressure $p^{\prime}$ and velocity ${\bf u}^{\prime}$ fields. The final stage of the algorithm comprises the correction step $p^{(n+1)}=p^{(n)}+p^{\prime}$ for calculating the pressure field, and the projection step ${\bf u}^{(n+1)}={\bf u}^{*}+{\bf u}^{\prime}$ for calculating the divergence-free velocity field.
In [16], the SIMPLE approach is extended to the immersed boundary framework by introducing a Lagarngian forces field ${\bf F}^{(n+1)}={\bf F}^{(n)}+{\bf F}^{\prime}$ and solving the following (semi-discretized) system of equations:
$\frac{3{\bf u}^{*}-4{\bf u}^{(n)}+{\bf u}^{(n-1)}}{2\Delta t}=-\nabla p^{(n)}+\frac{1}{Re}\nabla^{2}{\bf u}^{*}+R{\bf F}^{(n)}-\left({\bf u}\cdot\nabla{\bf u}\right)^{(n)}$ (8)
$\frac{3{\bf u}^{\prime}}{2\Delta t}=-\nabla p^{\prime}+R{\bf F}^{\prime}$ (9)
$R^{T}{\bf u}^{(n+1)}={\bf U}^{\Gamma}$ (10)
$\nabla\cdot{\bf u}^{(n+1)}=0$ (11)
where $R$ is the regulatizarion operator defined in Eq. (7), that smears the forces from the Lagrangian points to the adjacent Eulerian velocity locations by utilizing the discrete Dirac delta function proposed in [21], $R^{T}$ is the adjoint interpolation operator, defined in Eq. (6), that interpolates the intermediate velocity field onto the corresponding Lagrangian points, and ${\bf U}^{\Gamma}$ is the velocity value on the surface of the immersed body.

### v2

# A SIMPLE-Based Preconditioned Solver for the Direct-Forcing Immersed Boundary Method

@@ Abstract
We present a robust and scalable solver for direct-forcing immersed boundary simulations, based on a preconditioned SIMPLE algorithm. The method applies block elimination to the pressure-force coupled system, and utilizes the discrete Laplacian operator as an efficient preconditioner for the resulting Schur complement. We rigorously demonstrate the spectral equivalence between the Schur complement and the discrete Laplacian, ensuring convergence behavior that is independent of grid resolution and physical parameters. This enables accurate, stable, and efficient two-way coupled fluid-structure interaction (FSI) simulations with moving boundaries and significant added-mass effects. These simulations are all executable on standard computing platforms. Extensive validation and verification — including simulations of oscillating, sedimenting, and buoyant spheres, as well as configurations involving multiple immersed bodies — confirm the solver’s accuracy and efficiency across a broad range of FSI scenarios. The proposed approach introduces a novel and accessible framework for immersed boundary simulations requiring strong pressure–force coupling.
@@
Moving boundary simulations, Two way coupling FSI, Implicit immersed boundary method, Schur complement, Preconditioning
@@ 1 Introduction
Immersed boundary methods (IBMs) provide a versatile computational framework for fluid-structure interaction (FSI) problems involving complex or moving geometries, without requiring mesh regeneration or body-fitted grids. The coupled FSI problem is governed by fluid and structural equations, linked via interface constraints. These interface conditions enforce displacement and velocity continuity and stress balance, constituting a system of equations coupling velocity, pressure and Lagrangian forces. The numerical treatment of this system depends on the chosen solution strategy, broadly categorized as monolithic or partitioned. Monolithic methods solve all unknowns simultaneously, resulting in intrinsic coupling. In contrast, partitioned methods treat the fluid and structural problems within separate solver frameworks, exchanging information across the interface and enforcing coupling through explicit, semi-implicit, or fully implicit schemes. Explicit schemes apply fluid and structural updates sequentially within each time step. Semi-implicit methods introduce limited, typically one-way temporal coupling, updating one subsystem based on temporally lagged information from the other. Fully implicit schemes employ two-way coupling and iterate both subsystems to convergence at each time step. Explicit approaches can work well in steady or weakly coupled situations. However, they often fail when rapid transients occur, when the fluid and structure has comparable characteristic time scales, or when strong added-mass effects are present. In these cases, quasi-static assumptions no longer hold. To maintain stability and accurately conserve linear and angular momentum, one must use either a monolithic formulation or a partitioned scheme with semi- or fully-implicit coupling. Examples of such configurations include nearly neutrally buoyant particles [1], stiff elastic membranes [2], deformable capsules or vesicles modeled as nonlinear membranes [3], fluid–particle interaction during binder jet 3D printing [4], and fluid-structure piezoelectric interaction [5]. For a detailed overview of existing coupling schemes and their applications in incompressible FSI, see the comprehensive review [6].
Monolithic formulations offer strong coupling and numerical robustness, ensuring consistent enforcement of all constraints. These include algebraic preconditioning of the Newton iteration applied to large displacement FSI configurations [7], direct inversion of an extended Helmholtz operator via LU decomposition [8], Schur complement reduction with FFT-based Laplacian inversion under periodic boundary conditions [9, 10], Schur complement-based solution of the Stokes system accelerated by geometric multigrid preconditioning [11], and stream-function-vorticity formulations with efficiently precomputed fluid-structure coupling operators [12]. Despite their robustness, monolithic methods are often computationally expensive due to the need to solve large, indefinite, and ill-conditioned saddle-point systems. An additional drawback is that the numerical methodology must be developed essentially from scratch for each specific problem, which prevents reuse of legacy codes previously developed for fluid and structural solvers.
These limitations have contributed to the growing popularity of partitioned formulations over recent decades, which offer a practical alternative with greater modularity and scalability across diverse FSI applications. Partitioned formulations of IBMs, especially those based on the direct-forcing approach [13, 14], have shown particular effectiveness in incompressible flow simulations when employing semi-implicit or fully implicit coupling. These methods avoid assumptions about structural dynamics and instead enforce interfacial constraints directly through the momentum equations. The constraints are imposed by introducing Lagrangian force terms into the momentum equations to match the desired boundary velocities. Pressure and interface force fields act as distributed Lagrange multipliers (DLM), intrinsically coupled to each other and to the velocity field through the governing equations. Several works are worth mentioning in this context: in [15], it is proposed to combine all unknown Lagrangian forces into a single, fully coupled system to better reflect the parabolic nature of the Navier–Stokes equations. The works [16, 17] employ moving-least-squares reconstruction and isoparametric mapping to improve overall accuracy and robustness. The works [18, 19] aim to achieve high-order accuracy in enforcing the no-slip constraint, and in [20], Dirichlet and Robin boundary conditions are implemented with high accuracy in sharp corner geometries.
The immersed boundary projection method (IBPM), originally introduced in [21], has emerged as a particularly influential development within the family of partitioned direct-forcing immersed boundary formulations. It couples pressure and Lagrangian forces — both serving as Lagrange multipliers — within a unified projection framework and supports both semi-implicit and fully implicit [22] implementations. Over the past decade, IBPM has been effectively applied to simulate FSI involving rigid bodies, including sedimentation and particulate flows in two and three dimensions [23, 24, 25], as well as biologically inspired problems such as insect flight [26], vesicle dynamics over a wide range of Reynolds numbers [27, 28], and thermal convection involving heat transfer [29]. More recently, integration with SIMPLE [30] and PIMPLE [31, 32, 19] algorithms has considerably extended the method’s applicability, enabling robust simulations of stationary and arbitrarily moving geometries within widely used CFD platforms such as OpenFOAM.
While algorithmic advancements have broadened the applicability of immersed boundary methods, rigorous theoretical analysis remains limited, particularly for direct-forcing approaches. Most formal proofs of well-posedness and numerical stability have been established within the framework of variational formulations and finite element methods, typically tailored to fictitious domain methods, as demonstrated in the recent studies [33, 34] and references therein. Related efforts for continuous forcing formulation using finite-difference schemes have focused on proving the unconditional stability of backward Euler and Crank–Nicolson-like discretizations of the nonlinear immersed boundary terms [35], and on developing projection-based preconditioners to accelerate semi-implicit methods through Schur complement reduction [36]. Rigorous analysis of direct-forcing IBMs, particularly for methods based on finite volume or finite difference discretizations that use distributed Lagrange multipliers (DLMs) to enforce no-slip constraints is especially scarce. To the best of our knowledge, the only such effort is the preconditioned implicit direct-forcing (PIDF) method [37], which improves no-slip enforcement through iterative force computation. However, this method relies on empirically tuned parameters, which limits its generality.
The current work contributes to bridging this gap by introducing a new preconditioning strategy for a SIMPLE-based direct-forcing IBM. The method solves a regularized saddle-point system to simultaneously update pressure and Lagrangian force corrections, ensuring incompressibility and enforcement of the no-slip constraint. This saddle-point system is first reduced via a primal Schur complement which is in turn preconditioned by a Laplacian operator. Building on this formulation, we prove rigorously that the Laplacian is spectrally equivalent to the Schur complement, enabling an efficient and robust preconditioning strategy, and demonstrate numerically that the number of Krylov iterations remains low and independent of grid resolution and problem parameters. We demonstrate the applicability of the method to both moving boundary and two-way coupled FSI problems. First, the methodology is verified through simulations of flows induced by a transversely oscillating sphere. Second, the approach is challenged with more complex configurations involving multiple moving bodies, specifically porous spheres modeled as arrays of rigid sub-spheres. The simulations reveal distinct flow features associated with different array porosities, including variations in drag coefficients, phase shifts, and vortex evolution. Then we demonstrate the applicability of the approach to two-way FSI simulations, such as sedimenting and buoyant spherical particles. In all of the above mentioned configurations, the method shows good agreement with experimental and numerical reference data while preserving numerical stability and accuracy. Finally, we demonstrate that the method achieves close to linear and sub-linear scalability in computational time and memory usage, respectively, while maintaining notably low memory requirements, making it well-suited for realistic FSI simulations on standard workstations.
2 Theoretical background
2.1 Governing equations for the dynamics of a spherical particle in a Newtonian flow.
@@ 2 Theoretical background
In compliance with the direct-forcing IBM formulation, the incompressible flow around a solid body is governed by the non-dimensional Navier–Stokes (NS) and continuity equations, supplemented by the kinematic no-slip constraint on the body’s surface:
[…]
$\nabla\cdot{\bf u}=0$ $\forall\mathbf{x}\in\Omega,$ (2)
[…]
@@
Note that the last term of Eq. (1) constitutes the dynamic constraint, which defines the Lagrangian force density by which the immersed body acts on the surrounding flow, while Eq. (3) represents the kinematic no-slip constraint at the surface of the immersed body. If the kinematics of the immersed body is prescribed (e.g., the body moves with constant velocity or undergoes transverse oscillations), the configuration corresponds to a one-way coupled FSI problem. In this scenario, the system defined by Eqs. (1)–(3), complemented by appropriate boundary conditions, is mathematically closed and can be solved throughout the computational domain $\Omega$ [^1] Note that according to the direct-forcing IBM formulation, the flow is also computed inside the immersed body. However, for a rigid body, this internal flow field does not have physical significance.. In the case of two-way coupled FSI, the system of Eqs. (1)–(3) must be solved simultaneously with Newton’s laws governing the dynamics of the solid body. These non-dimensional equations, obtained by normalizing the moment-of-inertia tensor, time, and force density using $\rho_{f}D^{5}$, $D/U_{0}$, and $\rho_{f}U_{0}^{2}/D$, scales respectively, and formulated within the immersed boundary framework as suggested in [38], are written for a spherical particle as:
[…]
Here, the parameter $\theta=\rho_{p}/\rho_{f}$ is the ratio of particle to fluid density, and $Fr=gD/U_{0}^{2}$ is the Froude number. Importantly, for a spherical particle, the moment-of-inertia tensor reduces to a constant scalar, which allows Eqs. (4)–(5) to be conveniently formulated and solved in the inertial reference frame. However, for a general (non-spherical) particles, it is common to formulate the conservation of angular momentum in a body-attached reference frame. In this case, the coordinate system is typically aligned with the principal axes of the body, for which the moment-of-inertia tensor becomes a constant diagonal matrix. The angular velocity vector $\boldsymbol{\omega}_{cs}$ is then expressed in this body-fixed coordinate system. The transformation between the body-fixed and inertial frames is performed by utilizing rotation matrix which should be updated every time step as detailed in the next section.
2.2 Updating the rotation matrix.
The rotation matrix $\mathbf{R}^{(n+1)}$ is updated using the Rodriguez rotation formula based on the angular velocity vector. At each time step $n$, the angular velocity vector expressed in the body-fixed (co-rotating) reference frame, $\boldsymbol{\omega}_{cs}^{(n)}$, is first normalized to obtain the direction of the rotation axis:
$\hat{\boldsymbol{\omega}}_{cs}^{(n)}=\frac{\boldsymbol{\omega}_{cs}^{(n)}}{\|\boldsymbol{\omega}_{cs}^{(n)}\|}.$ (6)
The rotation angle $\gamma$ over the time interval $\Delta t$ is computed using a second-order Adams–Bashforth scheme:
$\gamma=\Delta t\left(\frac{3}{2}\|\boldsymbol{\omega}_{cs}^{(n)}\|-\frac{1}{2}\|\boldsymbol{\omega}_{cs}^{(n-1)}\|\right).$ (7)
A skew-symmetric matrix $\mathbf{W}$ is then constructed from the unit rotation axis $\hat{\boldsymbol{\omega}}_{cs}^{(n)}$:
$\mathbf{W}=\begin{bmatrix}0&-\hat{\omega}_{csz}^{(n)}&\hat{\omega}_{csy}^{(n)}\\ \hat{\omega}_{csz}^{(n)}&0&-\hat{\omega}_{csx}^{(n)}\\ -\hat{\omega}_{csy}^{(n)}&\hat{\omega}_{csx}^{(n)}&0\end{bmatrix}$, (8)
and the incremental rotation matrix is computed using the Rodrigues formula:
$\mathbf{R}_{\Delta}=I_{m}+\sin\gamma\,\mathbf{W}+(1-\cos\gamma)\,\mathbf{W}^{2},$ (9)
where $I_{m}$ is the identity matrix. Finally, the updated rotation matrix is obtained by composing the previous orientation $\mathbf{R}^{(n)}$ with the incremental rotation:
$\mathbf{R}^{(n+1)}=\mathbf{R}^{(n)}\mathbf{R}_{\Delta}.$ (10)
[…]
2.3 Interpolation and regularization operators.
Since the Eulerian and Lagrangian grids do not necessarily coincide, two adjoint operators are need to be introduced to facilitate the data transfer between them: the regularization operator $R$, which smears the Lagrangian forces onto the underlying Eulerian grid, and the interpolation operator $R^{T}$, which interpolates the Eulerian velocities to Lagrangian points. These operators are defined by utilizing the following discrete Dirac delta function, proposed in [39]:
@@
where $r$ represents the distance from the smeared point and $\Delta r$ represents the cell width in the $r$ direction. This discrete Dirac delta function gained popularity over the years [38, 40, 21, 30], due to its compact support of at most three discretization cells in each direction. [^2] Since the discrete Dirac delta function in Eq. (11) provides only a first order of spatial accuracy, other discrete delta functions can be also considered to improve the accuracy (with the price of less favorable sparsity), see [10]. The regularization and interpolation operators are defined using the discrete Dirac delta function from Eq. (11) as follows:
[…]
where $\mathbf{x}_{i}$ denotes the coordinates of the $i$-th velocity location, and $\mathbf{X}_{k}$ denotes the $k$-th Lagrangian point. These discrete summation formulas approximate the integrals under the assumption that the immersed surface has sufficiently smooth local curvature and that the Lagrangian points are uniformly distributed, with spacing comparable to that of the underlying uniform Eulerian grid. The notation $\delta^{3}$ refers to the three-dimensional discrete Dirac delta function. As long as the same delta function is used for both interpolation and regularization, the two operators are transposes of each other in matrix representation. For notational convenience, we denote the regularization operator by $R$ and the interpolation operator by $R^{T}$.
3 Method: time stepping and the pressure–force-density corrections system
3.1 One-way coupled moving boundary
We begin with a brief overview of the SIMPLE-IBM approach introduced in our recent work [30] by expressing the governing Eqs. (1)– (3) in a semi-discrete form:
$\frac{3{\bf u}^{*}-4{\bf u}^{(n)}+{\bf u}^{(n-1)}}{2\Delta t}=-\nabla p^{(n)}+\frac{1}{Re}\nabla^{2}{\bf u}^{*}+R{\bf F}^{(n)}-\left({\bf u}\cdot\nabla{\bf u}\right)^{(n)},$ (14)
$\frac{3{\bf u}^{\prime}}{2\Delta t}=-\nabla p^{\prime}+R{\bf F}^{\prime},$ (15)
$R^{T}{\bf u}^{(n+1)}={\bf U}^{\Gamma},$ (16)
$\nabla\cdot{\bf u}^{(n+1)}=0,$ (17)
where $R$ is the regularization operator defined in Eq. (13), which smears the forces from the Lagrangian points to the adjacent Eulerian velocity locations; $R^{T}$ is its adjoint interpolation operator, defined in Eq. (12), which interpolates the intermediate velocity field onto the corresponding Lagrangian points; and $\mathbf{U}^{\Gamma}$ is the prescribed velocity value on the surface of the immersed body. As already mentioned, in the case of one-way moving boundary coupling, the body kinematics is prescribed a priori (i.e., $\mathbf{x}^{n+1}$ and $\mathbf{U}^{\Gamma}$ are known from the kinematic law), and the above system is closed and can be solved subject to appropriate boundary conditions. A standard finite volume central-difference discretization was applied to all the spatial terms, while a second-order backward difference was utilized for discretization of all the temporal derivatives.